Vector v has a magnitude of 15 and a direction angle of 60 degrees. Calculate the resulting vector v + u, vector u has a magnitude of 25 and a direction angle of 140°
Answer by ikleyn(52778)

.

Vector v = (15*cos(60°),15*sin(60°)).


Vector u = (25*cos(140°),25*sin(140°)).


Vector u + v = (15*(cos(60°) + 25*cos(140°), 15*sin(60°) + 25*sin(140°))


Having clear instructions, you may complete calculations on your own.